OpenStudy (anonymous):

Solve on the interval [0,2π): (sinx+1)(2sin^2 x-3sinx - 2) = 0

OpenStudy (anonymous):

\[(a+1)(2a^2-3a-2)=0\] \[a=-1\]or\[ 2a^2-3a-2=0\] \[(2a+1)(a-2)\] \[a=-1/2,a=2\]

OpenStudy (anonymous):

A. x = 2π , x = π/2 , x = 5π/4 B. x = π , x = 2π/3 , x = 5π/3 C. x = 2π , x = π/2 , x = π/3 D. x = 3π/2 , x = 7π/6 , x = 11π/6

OpenStudy (anonymous):

put a=sin x

OpenStudy (anonymous):

From my work, it's D. Am I wrong?